for molecule HOOH (C2h symmetry), please find all vibrational modes with allowed fundamental transition.
Answer #1

To find the vibrational modes of a molecule HooH:

The given molecule is non linear for that fundamental vibration mode of molecule corrensponds transition from v=0 to=1.

For non linear molecule there will be 3N-6 number of vibrations.

Take HOOH by using formula 3N-6 here N =4 because there are four no of total atoms by substitute it form

=3×4-6

=12-6

=6

So there are six number of vibrational modes.

The normal modes of vibrations are asymmetric, symmetric, wagging,twisting and scissoring are shown figure 1CS Scanned with CamScanner

In 3N-6 it has 3N degrees of freedom including translation, rotation and vibration.

A non linear molecule can rotate about any of three mutual perpendicular axes and three rotational degrees of freedom

Finally the allowed fundamental transitions are shown figure レク CS Scanned with CamScanner